RunPlan 20170209 to March 6¶
Issues and tasks:¶
beam bump needs to be redone (Monday)
roof chiller intervention (Monday)
EMR daq stability
EMR spare board
AMI controller swap (Monday)
magnetic sweep need to be renewed
diffuser test @ 50A
Daily Preparations Schedules for Running.¶
Friday 10th:¶
Ramp DS to 200A.Fill SSU heliumCheck helium level monitor okayCheck trench water and target test, Henry Lock up hall ready for Saturday
Saturday 11th Expert Run:¶
beam bump tuning - Hayley (ISIS) - DONE
straight track data to test all systems - Need to restart and test EMR
Tracker startup from 9am - Mismatch in DAQ events to Trigger, Monday File Compactor script error detectionTarget tests from 10am
Network isolation test, disconnect optical fibre cable on main switch, isolating miceNet from external network (normal data taking activities should not be affected). 30 mins. - TBD
Sunday 12th No work planned¶
Monday 13th¶
Watchdog error fix
Left over from Saturday.
Tuesday 14th¶
Cooling channel magnet ramp
Wednesday 15th¶
Thursday 16¶
ramp of SSU (magnet team) - postponed
diffuser test (BLOC) - postponed
beam bump (Henry liaising with ISIS)
tracker calibration (Ed) - DONE
beam line tuning (Ao) - postponed
eventually straight tracks settings (alignment, calibration)
Friday 17¶
He refill in FC and SSD - DONE
beam bump - DONE (to be redone on Monday)
straight tracks: S5b
Sunday 18¶
10A test on SSU and SSD - DONE
TOF calibration: S6a, S6b
Wednesday 22¶
Tracker calibration
Reference run
Momentum scan with SSU on (205A), SSD on (205A), FDC off:
Cycle through H17a, H17b, H17c, H17d
Feb 15th to March 6th¶
Date
Feb 15
Feb 16
Feb 17
Feb 18
Feb 19
MOM
P. Franchini
P. Franchini
P. Franchini
P. Franchini
P. Franchini
BLOC
H. Nebrenski
T. Mohayai
T. Mohayai
T. Mohayai
T. Mohayai
SOC
D. Rajaram
D. Rajaram
D. Rajaram
D. Rajaram
D. Rajaram
Lead Shifter (00:00-08:00)
D. Maletic
D. Maletic
D.Maletic
D.Maletic
D.Maletic
Shifter (00:00-08:00)
D. Bowring
D. Bowring
D.Bowring
D.Bowring
D.Bowring
Lead shifter (08:00-16:00)
T. Mohayai
T. Mohayai
T. Mohayai
T. Mohayai
T. Mohayai
Shifter 2 (08:00-16:00)
A. DeBari
A. DeBari
A. DeBari
A. DeBari
A. DeBari
Lead shifter (16:00-00:00)
D. Neuffer
D. Neuffer
D.Neuffer
D.Neuffer
D.Neuffer
Shifter 2 (16:00-00:00)
A. Liu
A. Liu
A. Liu
A. Liu
A. Liu
Additional Personnel
Notes
Date
Feb 20
Feb 21
Feb 22
Feb 23
Feb 24
Feb 25
Feb 26
MOM
P. Franchini
P. Franchini
P. Franchini
P. Franchini
P. Franchini
P. Franchini
P. Franchini
BLOC
Y.Song
Y.Song
Y.Song
Y.Song
Y.Song
S. Boyd
S. Boyd
SOC
D. Rajaram
D. Rajaram
D. Rajaram
D. Rajaram
D. Rajaram
D. Rajaram
D. Rajaram
Lead Shifter (00:00-08:00)
Y.Song
Y.Song
Y.Song
Y.Song
Y.Song
E.Overton
E.Overton
Shifter (00:00-08:00)
C. Rogers
M Courthold
V. Bayliss
E.Overton
E.Overton
Lead Shifter(08:00-16:00)
D. Rajaram
D. Rajaram
D. Rajaram
D. Rajaram
D. Rajaram
S. Wilbur
S.Wilbur
shifter 2 (08:00-16:00)
H. Nebrenski
H. Nebrenski
H. Nebrenski
H. Nebrenski
H. Nebrenski
M.Federova
M.Federova
Lead shifter (16:00-00:00)
A. Dobbs
A. Dobbs
A. Dobbs
A. Dobbs
A. Dobbs
Y. Martyniak
Y. Martyniak
Shifter 2 (16:00-00:00)
K. Long
K. Long
K. Long
K. Long
K. Long
M.Savic
M. Savic
Additional Personnel
Notes
Date
Feb 27
Feb 28
Mar 1
Mar 2
Mar 3
Mar 4
Mar 5
Mar 6
MOM
M. Uchida
M. Uchida
M. Uchida
M. Uchida
M. Uchida
M. Uchida
M. Uchida
M. Uchida
BLOC
Y.Song
Y.Song
Y.Song
Y.Song
Y.Song
SOC
D. Rajaram
D. Rajaram
D. Rajaram
D. Rajaram
D. Rajaram
D. Rajaram
D. Rajaram
D. Rajaram
Lead Shifter (00:00-08:00)
E.Overton
V. Bayliss
C. Rogers
Shifter (00:00-08:00)
M.Bogomilov
M.Bogomilov
M.Bogomilov
M.Bogomilov
M.Bogomilov
Lead Shifter(08:00-16:00)
S. Wilbur
S. Wilbur
S.Wilbur
A. Blondel
A. Blondel
A. Blondel
A. Blondel
A. Blondel
shifter 2 (08:00-16:00)
M.Federova
M.Federova
M.Federova
P. Kyberd
P. Kyberd
S. Boyd
S. Boyd
S. Boyd
Lead shifter (16:00-00:00)
J. Martyniak
D. Rajaram
D. Rajaram
Y.Song
Y.Song
Y.Song
Y.Song
Y.Song
Shifter 2 (16:00-00:00)
M.Savic
M.Savic
M.Savic
A. Bross
A. Bross
A. Bross
A. Bross
A. Bross
Additional Personnel
Notes
Data to take:¶
Field On Data¶
Take all H13a then all H14a-c then H13b then H13c (so that diffuser settings are used last).
Task H13 - Emittance Reduction¶
Task
Momentum
# Particle Triggers
Beamline Tag
CoolingChannel Tag
Similar runs
Final Event Tally (k)
H13a
140
630K TOF2
3-140+M3-Test2
2016-05-4
Task H14 - Magnet alignment¶
Task
Momentum
# Particle Triggers
Beamline Tag
CoolingChannel Tag
Similar runs
Final Event Tally (k)
H14a
170
1.5e5 TOF1
3-170+M3-Test1
2016-05-4
H14b
200
1.5e5 TOF1
3-200+M3-Test1
2016-05-4
H14c
240
1.5e5 TOF1
3-240+M3-Test1
2016-05-4
Task H13 - Emittance Reduction with diffuser (after H13a and H14 are completed)¶
Task
Momentum
# Particle Triggers
Beamline Tag
CoolingChannel Tag
Similar runs
Final Event Tally (k)
H13b
140
315K TOF2
6-140+M3-Test2
2016-05-4
H13c
140
200K TOF2
10-140+M3-Test3
2016-05-4
Data to take after changing the CC settings:¶
Task H15 - Emittance Reduction¶
Task
Momentum
# Tof2 Triggers
Beamline Tag
CoolingChannel Tag
Similar runs
Final Event Tally (k)
H15a
140
630K TOF2
3-140+M3-Test2
2016-05-1
H15b
140
315K TOF2
6-140+M3-Test2
2016-05-1
H15c
140
200K TOF2
10-140+M3-Test3
2016-05-1
Task H16 - Magnet alignment¶
Task
Momentum
# Particles Triggers
Beamline Tag
CoolingChannel Tag
Similar runs
Final Event Tally (k)
H16a
170
5e5 TOF1
3-170+M3-Test1
H16b
200
5e5 TOF1
3-200+M3-Test1
H16c
240
5e5 TOF1
3-240+M3-Test1
Straight track data (fallback)¶
Detector alignment¶
Task
Momentum
# Particle Triggers
Beamline Tag
ColingChannel Tag
Similar runs
Final Event Tally (k) @ 10am
S5b
300
1000k at TOF1
300MeV+pi pa82
StepIV-Off
8030
400
S5c
400
1000k at TOF1
400MeV+pi pa82
StepIV-Off
8041
TOF calibration¶
Task
Momentum
# Particle Triggers
Beamline Tag
CoolingChannel Tag
Final Event Tally (k) @ 10am
S6a
150
2M @ TOF1
ToF-Calib+150
StepIV-Off
448
S6b
300
1M @ TOF1
ToF-Calib+300
StepIV-Off
220
Step4_Pionic_beam_Tags
Files (1)
Updated by Franchini, Paolo almost 7 years ago
ยท 51 revisions